what is the value of log 1.8 ??? tell the method

Answer by anantha(86)

sol:
i.8 can be written as 18/10
now log 1.8=log (18/10)
log (18/10)=log (2*9/2*5)
=log(9/5)
=log (a/b)=loga-logb
log(9/5)=log 9-log 5
=log(3*3)-log 5
=log(3^2)-log 5
=2log3-log5 (i.e;log m^n=n logm)
log 1.8=2 log3-log 5